Day operations


  • Mounted the unit on the telescope
    • REMAIND: formal request to LBTO about mount the Swagelock of the telescope cooling lines on the swing-arm interface
  • Removed the mirror cover (M2 retracted)
  • gap has no contamination: TSS enabled (switch in the swing-arm cabinet)
  • at zenith tested set_and_flat: OK


  • Dome opened at about 01:40UT
  • No mirror cover: Software started to handle TSS (checked communication between AdSec Arbitrator and AOS for windspeed data)
  • M2 is retracted: no setting of the shell is allowed during the night
  • Temperature cooling oscillation:
    • there is about 1.5C PtV oscillation, about 100s period
    • The oscillation is not always present. See for instance from 01:30UT to 04:00UT
    • The temperature variation induces a correlated oscillation on the reading of capacitive sensor (possibly because of mechanical deformation)

-- ArmandoRiccardi - 19 Sep 2010
Topic revision: r3 - 20 Sep 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback